# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### v40

@@ 3 General Model
As in Section 2, there are two risk-neutral agents, the insider and the market maker. At $t=0$, the insider observes a signal that informs him of the probability distribution over the set of possible $t=1$ states, $X=\{x_{1},x_{2},\cdots\}$ (which is countably infinite, with finite $X$ included as a special case). The market maker has a Bayesian prior over the possible signals. At $t=0$, there is a complete market of AD securities for $t=1$ states. [^11] The model can be adapted to the incomplete market case by partitioning the state space. In this paper, we focus on the complete market case. There is a risk-free asset in perfectly elastic supply at risk-free rate zero.
@@
The set of possible signals is $S=\{s_{1},\cdots,s_{K}\}$, with the market maker holding prior $\pi_{0}(s_{k})$, $1\leq k\leq K$. Conditional on a signal $s_{k}$, the probabilities of $t=1$ states are given by $\eta(\,\cdot\,|s_{k})\colon X\rightarrow[0,\infty)$. After observing $s_{k}$, the insider chooses a portfolio $W(\,\cdot\,|s_{k})\colon X\rightarrow\mathbb{R}$ where $W(x_{i}|s_{k})$ is his order for security $x_{i}$ conditional on $s_{k}$.
@@
The market maker prices contracts conditional on $\omega=(\omega_{i})$ exactly as in the AD formulation. Accordingly, all results derived in the AD formulation translate directly to options.
@@ 4 Cross-Market Inference
The market maker’s inference problem is infinite-dimensional: as noisy order flow arrives across many claims, he must infer the informed portfolio behind it, which is unrestricted. Making the underlying economic intuition precise therefore requires going beyond finite-dimensional Bayesian updating. Once we do, the takeaway is simple. His sufficient statistic is the degree to which order flow aligns with the insider-demand portfolio he expects under each signal—a direct generalization of the cross-market imbalance $\Delta(\omega)$ of Section 2.
@@ 4.1 The Posterior and Pricing Kernel
The market maker treats $s_{k}$ as the latent parameter, with prior $\pi_{0}(s_{k})$ for $k=1,\ldots,K$, and updates from the realized order flow $\omega$. Let $\mathbb{P}_{\scriptscriptstyle W(\,\cdot\,|s_{k})}$ denote the likelihood of $\omega$ under (8) when the insider’s mean order is $W(\,\cdot\,|s_{k})$, and let $\mathbb{P}_{\!\scriptscriptstyle 0}$ denote the noise-only likelihood (i.e., $W\equiv 0$). Since the noise-only likelihood does not depend on signal, the market maker can summarize the likelihood of $\omega$ conditional on $s_{k}$ via the likelihood ratio [^13] Formally, $\mathbb{P}_{\scriptscriptstyle W(\,\cdot\,|s_{k})}$ and $\mathbb{P}_{\!\scriptscriptstyle 0}$ are probability measures on $\Omega$. The likelihood ratio (9) is their Radon-Nikodym derivative.
@@ 5.1 First-Order Condition
The general price-impact term $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)$ is written for an arbitrary trade package $v$. The primitive empirical object is the cross-impact kernel—how signed order flow in one claim moves the price of another. Corollary 5.2 below isolates this kernel by specializing to the case where $v$ consists of one share of claim $x_{j}$. In that case, its impact exposure collapses to $\Pi_{insider}(v,\,\cdot\,;\widetilde{W})=\widetilde{W}(x_{j}\mid\,\cdot\,)/\sigma_{j}^{2}$—that is, how much a marginal trade in $x_{j}$ tilts the market maker’s posterior toward a given signal is exactly how much $x_{j}$ the market maker expects the insider to demand under that signal, adjusted for noise.
@@
Under the market maker’s belief $\widetilde{W}$, the price impact of marginal order flow for $x_{j}$ on the price of another claim $x_{i}$ given realized aggregate order flow $\omega$—is
@@
Averaging (20) over the distribution of realized order flow generated by portfolio $W$ gives the expected cross impact
@@
One can read $\widetilde{W}(x_{j}\mid\,\cdot\,)$ as the market maker’s story-by-story “informed-demand template” for $x_{j}$ and $\eta(x_{i}\mid\,\cdot\,)$ as $x_{i}$’s payoff profile across the same stories.
@@ 5.2 No-Arbitrage
[…]
$\Pi_{insider}(W,\,\cdot\,;\widetilde{W})=0,$
i.e., $W$ is orthogonal to the market maker’s belief about informed demand. These are “stealth” portfolios—they sit in the null space of the cross-impact kernel, so scaling them does not tilt the market maker’s posterior and therefore does not move prices. If such a zero-impact portfolio yielded nonzero payoff to the insider, the insider could lever it up without moving prices, generating an arbitrarily large payoff. This is Theorem 5.3.
@@ 6 Equilibrium
Up to this point we have solved each side of the market taking the other as given: the market maker’s pricing kernel as a mapping from cross-market order flow to the cross section of claim prices (Section 4), and the insider’s optimal portfolio in the same rich claim space (Section 5). In equilibrium, the conjectured “informed-demand template” that drives pricing must coincide with the insider’s optimal trades under that very pricing kernel—a joint fixed point of two high-dimensional problems. A key insight is that the trading game is isomorphic to a canonical game that is invariant to the information structure and noise intensities. [^15] This canonical-game reduction is fully robust. It holds whether the signal and state spaces are discrete or continuous. We keep the exposition in the finite-$S$ case here. Fully rigorous arguments for the continuous case are provided in Keller and Tseng (2026). This sets up a striking reduction—equilibrium, an infinite-dimensional fixed point a priori, is characterized by a single endogenous constant.
@@
For expositional convenience, we maintain Assumption 1 in this section. It is without loss of generality and serves only to streamline the derivations. Starting from any prior $\pi_{0}$, we can rescale payoffs and work as if the prior were uniform on $S$ by replacing $\eta(\,\cdot\,|s_{k})$ with $\frac{\pi_{0}(s_{k})}{1/|S|}\,\eta(\,\cdot\,|s_{k})$ for each $s_{k}\in S$. Once we characterize equilibrium under the uniform prior, comparative statics for arbitrary priors follow immediately. For instance, if two securities both yield higher payoffs under some $s_{k}$ compared to the other signals, then their cross price impact increases with the prior probability of $s_{k}$. For instance, if two securities tend to pay off together under some $s_{k}$, then putting more prior weight on $s_{k}$ makes their co-movement in order flow more diagnostic, so their equilibrium cross price impact increases with $\pi_{0}(s_{k})$. This generalizes from the toy model of Section 2; see footnote LABEL:fn:_general_prior.
@@ 6.1 Canonical Game
The information intensity matrix $\bf L$} is the cross-market analogue of the information intensity parameter $\sigma_{v}/\sigma_{\varepsilon}$ from the single-asset Kyle setting (6). It aggregates the (noise-adjusted) co-variations of the signal-implied payoff profiles $\{\eta(\,\cdot\,\mid s_{k})\}_{k}$ across claims, so a larger $\bf L$} means that cross-market order flow is more diagnostic of which information story is active. [^16] Here, “larger” is in the positive-semidefinite order. This order is partial—unlike the single-asset setting, information intensity is not always directly comparable between two specifications.
@@
The information-unit conversion (22) strips away the idiosyncrasies of the payoff distributions and noise intensities and isolates the strategic core of the trading problem. In information units, the payoff-relevant span becomes orthonormal—each trading direction carries one unit of signal relative to noise. The trading problem can then be read as trading orthogonal information contracts—one per information story—under unit noise intensity. The next theorem records this equivalence.
@@ 6.2 One-Parameter Equilibrium
The canonical-game reduction sets up the central observation. Despite the trading game being infinite-dimensional, its equilibrium restriction is extremely tight—it collapses to a single endogenous scalar.
@@
Imposing insider optimality within this symmetric class reduces the fixed point condition to a single scalar equation in $\alpha$. Theorem 6.3 shows that it admits a solution $\alpha^{*}>0$, delivering equilibrium strategy matrix $D^{*}=\alpha^{*}{\bf Q}$. Thus, $\alpha^{*}$ is the sole endogenous constant characterizing equilibrium.
@@
The next subsection demonstrates the especially transparent binary-signal case (which nests the toy setting of Section 2), where inference collapses to a single log-odds statistic.
@@ 6.3 Example - Binary Signal Case
Thus observing order flow is equivalent to observing the “evidence score” $Z$: large $Z$ is strong evidence for $s_{1}$, while large negative $Z$ favors $s_{2}$. In the toy setting of Section 2, the evidence score reduces to the normalized cross-market imbalance $\Delta(\omega)$ in (2), with $Z=2\alpha\,\Delta(\omega)$. Accordingly, in the volatility interpretation of Section 2, large $Z$ is strong evidence for high volatility.
@@
where the expectations are taken under (27) (and hence are functions of $\alpha$). [^19] This is an instance of the general equilibrium equation (A.23). Here $1-\mathbb{E}[p_{1}]$ is the average “wedge” the insider earns from the active story, while $p_{1}p_{2}=p_{1}(1-p_{1})$ is the posterior variance: it is largest when $Z\approx 0$ (the market is unsure) and small when $|Z|$ is large (the market is confident), making precise how marginal scaling of the spread further sharpens inference and erodes marginal information rents.
The equilibrium scale $\alpha^{*}$ solves $\Phi(\alpha)=0$, i.e., the marginal gain from increasing exposure is offset by the marginal information leakage. We have $\Phi(0)=\tfrac{1}{2}>0$, while $\Phi(\alpha)<0$ for sufficiently large $\alpha$, and moreover $\Phi(\alpha)\uparrow 0$ as $\alpha\to\infty$. [^20] The limit $\Phi(\alpha)\uparrow 0$ follows from Fatou’s Lemma. Existence of $\alpha^{*}>0$ with $\Phi(\alpha^{*})=0$ follows by the Intermediate Value Theorem. Hence there exists $\alpha^{*}>0$ solving $\Phi(\alpha)=0$, which proves Theorem 6.3 in this binary signal case.
@@ 6.4 Comparative Statics of $\alpha^{*}$
As $K$ increases, there are more distinct payoff-distribution “stories” the market must disentangle from cross-contract order flow, so equilibrium informed trading becomes more aggressive. At the same time, a larger $K$ also increases how differently informed demand loads across stories, making order flow more diagnostic and steepening the informational (price-impact) cost of further scaling. The net effect is that $\alpha^{*}(K)$ rises with $K$ but with diminishing increments: $\alpha^{*}(K)$ is increasing and concave.
@@
$i,j$ Indices for Arrow-Debreu states/claims (e.g., option strikes in the options implementation via (7).)
@@
$\omega=(\omega_{i})_{i=1}^{N}$ Aggregate order flow across claims received by the market maker; $\omega_{i}=W^{*}(x_{i}\mid s_{k})+\varepsilon_{i}$.
@@ 7 Price Discovery
$\begin{array}[]{c}\text{\scriptsize signal}\\[-0.86108pt] s_{k}\end{array}\;\raisebox{-6.88889pt}{$\Rightarrow$}\;\begin{array}[]{c}\text{\scriptsize informed demand}\\[-0.86108pt] W^{*}(\,\cdot\,\mid s_{k})\end{array}\;\raisebox{-6.88889pt}{$\Rightarrow$}\;\begin{array}[]{c}\text{\scriptsize order flow}\\[-0.86108pt] \omega\end{array}\;\raisebox{-6.88889pt}{$\Rightarrow$}\;\begin{array}[]{c}\text{\scriptsize prices}\\[-0.86108pt] P^{*}(\,\cdot\,\mid\omega)\end{array}\;\raisebox{-6.88889pt}{$\Rightarrow$}\;\begin{array}[]{c}\text{\scriptsize price impact}\\[-0.86108pt] \Lambda\end{array}$ (29)
The private signal $s_{k}$ induces equilibrium informed demand $W^{*}(\,\cdot\,\mid s_{k})$, which is part of aggregate order flow $\omega$. Conditioning on $\omega$, the market maker sets Arrow-Debreu state prices $P^{*}(\,\cdot\,\mid\omega)$; the within- and cross-price responses are collected in the price impact matrix $\Lambda$. The standard formula (7) translates these Arrow–Debreu quantities to options.
@@ 7.1 Informed Demand
The expression (30) can be read as a simple three-step recipe for constructing the insider’s portfolio: [^21] The informed demand of Section 2 is the orthogonal-payoff special case of this general intuition: because payoff profiles do not overlap, the liquidity adjustment in ${\bf L}^{-1}$ is mute, so the insider’s problem reduces to choosing the long-short “information trade” and a single equilibrium scale $\alpha^{*}$.
@@
  • Step 2 (Liquidity adjustment). Rotate and scale the signal tilt by the matrix ${\bf L}^{-1}$. $\bf L$} summarizes how informative order flow is to market makers as they observe joint trading across contracts; applying ${\bf L}^{-1}$ therefore damps positions in directions where order flow is more revealing and hence more costly in terms of price impact.
@@
Left column: the insider’s private signal, indicated by the shaded distributions. Right column: the corresponding informed demand $W^{*}$.
@@
Suppose the signal selects between two normal payoff distributions with a common variance and means $\mu_{1}>\mu_{2}$. Private information therefore concerns the level (expected payoff) of the underlying, as in Kyle (1985), but the insider can trade options. In terms of the information flow (29), the equilibrium maps this mean-shift signal into a standard vertical spread option portfolio. [^22] See Hull (2015) for standard definitions of vertical spreads.
@@
Portfolio intuition. A high-mean signal $s_{1}$ induces a bull spread (panel 3b): a long–short vertical position that profits from upward revaluation while limiting exposure in far out-of-the-money states. Conversely, a low-mean signal $s_{2}$ induces a bear spread (panel 3d), concentrating downside exposure while economizing on extreme strikes that contribute little to expected profits under the signal-implied distribution.
@@
Left column: the insider’s private signal, indicated by the shaded payoff distributions. Right column: the corresponding informed demand $W^{*}$.
@@
Price Impact Three states are indicated: $x_{1}$, $x_{2}$, and $x_{3}$.
@@
Suppose the signal selects between two normal payoff distributions with a common mean and variances $\sigma_{1}^{2}>\sigma_{2}^{2}$. Private information therefore concerns the volatility of the payoff rather than its level. In terms of the information flow (29), equilibrium maps this volatility signal into the same workhorse option portfolios used throughout practice: a straddle when volatility is high, and a butterfly when volatility is low. [^23] See again Hull (2015) for standard definitions of straddles and butterflies.
@@
Portfolio intuition. Under the high-volatility signal $s_{1}$, the distribution places relatively more weight on tail outcomes. Accordingly, the informed demand loads on the wings and avoids the center (panel 4b), producing a straddle-type position—a wing-loaded long-volatility portfolio. Conversely, under the low-volatility signal $s_{2}$ the payoff is more concentrated near the mean. Panel 4d shows that informed demand flips sign across strikes, producing a butterfly position: it concentrates exposure near the center and shorts wing payoffs that are unlikely to realize under low volatility.
@@
Left column: signal-implied payoff distributions (shaded). Right column: corresponding informed demand $W^{*}$.
@@
Suppose the signal selects between two skew-normal payoff distributions that share the same mean and variance but differ in their skewness parameters, with $\alpha_{1}>\alpha_{2}$. Private information therefore concerns asymmetry—how probability mass shifts between left- and right-tail outcomes holding the first two moments fixed. In equilibrium, this skewness signal maps into a standard skew trade: an option portfolio that tilts exposure toward the tail favored by the signal while offsetting the opposite tail.

### v41

@@ 3 General Model
As in Section 2, there are two risk-neutral agents, the insider and the market maker. At $t=0$, the insider observes a signal that reveals the probability distribution over the set of possible $t=1$ states, $X=\{x_{1},x_{2},\cdots\}$ (which is countably infinite, with finite $X$ included as a special case). The market maker has a Bayesian prior over the possible signals. At $t=0$, there is a complete market of AD securities for $t=1$ states. [^9] The model can be adapted to the incomplete market case by partitioning the state space. In this paper, we focus on the complete market case. There is a risk-free asset in perfectly elastic supply at a risk-free rate of zero.
@@
The set of possible signals is $S=\{s_{1},\cdots,s_{K}\}$, with prior probabilities $\pi_{0}(s_{k})$, $1\leq k\leq K$. Conditional on a signal $s_{k}$, the probabilities of $t=1$ states are given by $\eta(\,\cdot\,|s_{k})\colon X\rightarrow[0,\infty)$. After observing $s_{k}$, the insider chooses a portfolio $W(\,\cdot\,|s_{k})\colon X\rightarrow\mathbb{R}$ where $W(x_{i}|s_{k})$ is his order for security $x_{i}$ conditional on $s_{k}$.
@@
The market maker prices contracts conditional on $\omega=(\omega_{i})$, and the rest of the model proceeds exactly as in the AD formulation. Conversely, all results derived in the AD formulation specialize immediately to the options setting.
@@ 4 Cross-Market Inference
The market maker’s inference problem is infinite-dimensional: as noisy order flow arrives across many claims, he must infer the informed portfolio behind it, which is unrestricted. Making the underlying economic intuition precise therefore requires going beyond the textbook finite-dimensional Bayesian formalism. Once we do, the takeaway is simple. His sufficient statistic is the degree to which order flow aligns with the insider-demand portfolio he expects under each signal—a direct generalization of the cross-market imbalance $\Delta(\omega)$ of Section 2.
@@ 4.1 The Posterior and Pricing Kernel
The market maker treats $s_{k}$ as the latent parameter, with prior $\pi_{0}(s_{k})$ for $k=1,\ldots,K$, and updates his belief based on the realized order flow $\omega$. Let $\mathbb{P}_{\scriptscriptstyle W(\,\cdot\,|s_{k})}$ denote the likelihood of $\omega$ under (8) when the insider’s mean order is $W(\,\cdot\,|s_{k})$, and let $\mathbb{P}_{\!\scriptscriptstyle 0}$ denote the noise-only likelihood (i.e., $W\equiv 0$). Since the noise-only likelihood does not depend on signal, the market maker can summarize the likelihood of $\omega$ conditional on $s_{k}$ via the likelihood ratio [^11] Formally, $\mathbb{P}_{\scriptscriptstyle W(\,\cdot\,|s_{k})}$ and $\mathbb{P}_{\!\scriptscriptstyle 0}$ are probability measures on $\Omega$. The likelihood ratio (9) is their Radon-Nikodym derivative.
@@ 5.1 First-Order Condition
The general price-impact term $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)$ is written for an arbitrary trade package $v$. The primitive empirical object is the cross-impact kernel—how signed order flow in one claim moves the price of another. When $v$ consists of one share of claim $x_{j}$, its impact exposure collapses to $\Pi_{insider}(v,\,\cdot\,;\widetilde{W})=\widetilde{W}(x_{j}\mid\,\cdot\,)/\sigma_{j}^{2}$. In words, order flow in $x_{j}$ pushes beliefs toward whichever story predicts heavier insider demand for $x_{j}$, adjusted for noise. Corollary 5.2 characterizes the resulting cross-impact kernel.
@@
Under the market maker’s belief $\widetilde{W}$, the price impact of marginal order flow for $x_{j}$ on the price of another claim $x_{i}$, given realized aggregate order flow $\omega$, is
@@
Averaging (20) over the distribution of realized order flow gives the expected cross impact
@@
$\widetilde{W}(x_{j}\mid\,\cdot\,)$ is the market maker’s story-by-story predicted informed demand for $x_{j}$, and $\eta(x_{i}\mid\,\cdot\,)$ is $x_{i}$’s corresponding payoff profile across those same stories.
@@ 5.2 No-Arbitrage
No-arbitrage in our setting admits an intuitive characterization: any portfolio that generates no price impact cannot yield a positive payoff for the insider. Given market maker belief $\widetilde{W}$, we call a portfolio $W$ a zero–price-impact portfolio if its impact exposure (14) vanishes,
$\Pi_{insider}(W,\,\cdot\,;\widetilde{W})=0.$
These are “stealth” portfolios—they look like noise trades to the market maker (i.e., they are orthogonal to the market maker’s belief about informed demand), so scaling them does not tilt the market maker’s posterior and therefore does not move prices. If such a zero-impact portfolio yielded a nonzero payoff to the insider, the insider could lever it up without moving prices, generating an arbitrarily large payoff. This is Theorem 5.3.
@@ 6 Equilibrium
Up to this point we have solved each side of the market taking the other as given: the market maker’s pricing kernel as a mapping from cross-market order flow to the cross-section of claim prices (Section 4), and the insider’s optimal portfolio in the same rich claim space (Section 5). In equilibrium, the conjectured “informed-demand template” that drives pricing must coincide with the insider’s optimal trades under that very pricing kernel—a joint fixed point of two high-dimensional problems. A key insight is that the trading game is isomorphic to a canonical game that is invariant to the information structure and noise intensities. This sets up a striking reduction—equilibrium, an infinite-dimensional fixed point a priori, is characterized by a single endogenous constant. (The economic intuition underlying this reduction is robust whether the signal and state spaces are discrete or continuous; a formal treatment of the continuous case can be found in Keller and Tseng (2026).)
@@
For expositional convenience, we maintain Assumption 1 in this section. This is without loss of generality and serves only to streamline the derivations. Starting from any prior $\pi_{0}$, we can rescale payoffs and work as if the prior were uniform on $S$ by replacing $\eta(\,\cdot\,|s_{k})$ with $\frac{\pi_{0}(s_{k})}{1/|S|}\,\eta(\,\cdot\,|s_{k})$ for each $s_{k}\in S$. Once we characterize equilibrium under the uniform prior, comparative statics for arbitrary priors follow immediately. For instance, if two securities tend to pay off together under some $s_{k}$, then putting more prior weight on $s_{k}$ makes their co-movement in order flow more diagnostic, so their equilibrium cross price impact increases with $\pi_{0}(s_{k})$. This generalizes from the toy model of Section 2; see footnote 7.
@@ 6.1 Canonical Game
The information intensity matrix $\bf L$} is the cross-market analogue of the information intensity parameter $\sigma_{v}/\sigma_{\varepsilon}$ from the single-asset Kyle setting (6). It aggregates the (noise-adjusted) covariations of the signal-implied payoff profiles $\{\eta(\,\cdot\,\mid s_{k})\}_{k}$ across claims, so a larger $\bf L$} (in the natural matrix order) means that cross-market order flow is more diagnostic of which information story is active.
@@
The information-unit conversion (22) strips away the idiosyncrasies of the payoff distributions and noise intensities and isolates the strategic core of the trading game. In information units, the payoff-relevant span becomes whitened (orthonormal)—each trading direction carries one unit of signal relative to noise. The trading game can then be read as trading orthogonal information contracts—one per information story—under unit noise intensity. The next theorem records this equivalence.
@@ 6.2 One-Parameter Equilibrium
We now close the model via the canonical-game reduction, which reveals the extremely tight equilibrium restriction of the general trading game.
@@
Imposing insider optimality within this symmetric class reduces the fixed point condition to a single scalar equation in $\alpha$. Theorem 6.3 shows that it admits a solution $\alpha^{*}>0$, with corresponding equilibrium strategy matrix $D^{*}=\alpha^{*}{\bf Q}$. Thus, $\alpha^{*}$ is the sole endogenous constant characterizing equilibrium.
@@
The next subsection demonstrates the especially transparent binary-signal case (which nests the setting of Section 2), where inference collapses to a single log-odds statistic.
@@ 6.3 Example - Binary Signal Case
Thus, observing order flow is equivalent to observing the “evidence score” $Z$: large $Z$ is strong evidence for $s_{1}$, while large negative $Z$ favors $s_{2}$. In the toy setting of Section 2, the evidence score reduces to the normalized cross-market imbalance $\Delta(\omega)$ in (2), with $Z=2\alpha\,\Delta(\omega)$. Accordingly, in the volatility interpretation of Section 2, large $Z$ is strong evidence for high volatility.
@@
Here $1-\mathbb{E}[p_{1}]$ is the insider’s average rent wedge from the active story, while $p_{1}p_{2}=p_{1}(1-p_{1})$ is the market’s residual uncertainty (the posterior variance of the binary signal). The residual uncertainty is maximized when $Z\approx 0$ (the market is unsure) and it collapses as $|Z|$ grows (the market becomes confident). Thus, a marginal increase in the trading scale $\alpha$ reveals the most information precisely when the market is unsure, sharpening inference and compressing marginal information rents.
The equilibrium scale $\alpha^{*}$ solves $\Phi(\alpha)=0$, i.e., the marginal gain from increasing exposure is offset by the marginal information leakage. We have $\Phi(0)=\tfrac{1}{2}>0$, while $\Phi(\alpha)<0$ for sufficiently large $\alpha$, and moreover $\Phi(\alpha)\uparrow 0$ as $\alpha\to\infty$. [^15] The limit $\Phi(\alpha)\uparrow 0$ follows from Fatou’s Lemma. Existence of $\alpha^{*}>0$ with $\Phi(\alpha^{*})=0$ follows by the Intermediate Value Theorem. Hence, there exists $\alpha^{*}>0$ solving $\Phi(\alpha)=0$, which proves Theorem 6.3 in this binary signal case.
@@ 6.4 Comparative Statics of $\alpha^{*}$
As $K$ increases, there are more distinct payoff-distribution “stories” the market must disentangle from cross-contract order flow, so equilibrium informed trading becomes more aggressive. At the same time, a larger $K$ also increases the extent to which informed demand differs across stories, making order flow more diagnostic and steepening the informational (price-impact) cost of further scaling. The net effect is that $\alpha^{*}(K)$ rises with $K$ but with diminishing increments: $\alpha^{*}(K)$ is increasing and concave.
@@
$i,j$ Indices for Arrow-Debreu states/claims (e.g., option strikes in the options implementation via (7)).
@@
$\omega=(\omega_{i})_{i\geq 1}$ Aggregate order flow across claims received by the market maker; $\omega_{i}=W^{*}(x_{i}\mid s_{k})+\varepsilon_{i}$.
@@ 7 Price Discovery
$\begin{array}[]{c}\text{\scriptsize signal}\\[-0.86108pt] s_{k}\end{array}\;\raisebox{-6.88889pt}{$\Rightarrow$}\;\begin{array}[]{c}\text{\scriptsize informed demand}\\[-0.86108pt] W^{*}(\,\cdot\,\mid s_{k})\end{array}\;\raisebox{-6.88889pt}{$\Rightarrow$}\;\begin{array}[]{c}\text{\scriptsize order flow}\\[-0.86108pt] \omega\end{array}\;\raisebox{-6.88889pt}{$\Rightarrow$}\;\begin{array}[]{c}\text{\scriptsize pricing kernel}\\[-0.86108pt] P^{*}(\,\cdot\,\mid\omega)\end{array}\;\raisebox{-6.88889pt}{$\Rightarrow$}\;\begin{array}[]{c}\text{\scriptsize price impact}\\[-0.86108pt] \Lambda\end{array}$ (29)
The private signal $s_{k}$ induces equilibrium informed demand $W^{*}(\,\cdot\,\mid s_{k})$, which is part of aggregate order flow $\omega$. Conditioning on $\omega$, the market maker sets Arrow-Debreu state prices $P^{*}(\,\cdot\,\mid\omega)$; the within- and cross-price responses are collected in the price impact matrix $\Lambda$. The standard formula (7) translates these Arrow-Debreu quantities to options.
@@ 7.1 Informed Demand
The expression (30) can be read as a simple three-step recipe for constructing the insider’s portfolio:
@@
  • Step 2 (Liquidity adjustment). Rotate and scale the signal tilt by the matrix ${\bf L}^{-1}$. $\bf L$} summarizes how informative order flow is to market makers. Applying ${\bf L}^{-1}$ therefore reduces positions in directions where order flow is more revealing and hence more costly in terms of price impact.
@@
The informed demand of Section 2 is the orthogonal-payoff special case of this general intuition: because payoff profiles do not overlap, the liquidity adjustment ${\bf L}^{-1}$ is inactive, so the insider’s portfolio problem reduces to choosing the long-short information trade and the equilibrium scale $\alpha^{*}$.
@@
Left column: the insider’s private signal, indicated by the shaded distributions. Right column: the corresponding informed demands $W^{*}$.
@@
Suppose the signal selects between two normal payoff distributions with a common variance and means $\mu_{1}>\mu_{2}$. Private information therefore concerns the level (expected payoff) of the underlying, as in Kyle (1985), but the insider can trade options. In terms of the information flow (29), equilibrium maps this mean-shift signal into a standard vertical spread option portfolio. [^16] See Hull (2015) for standard definitions of vertical spreads.
@@
Portfolio intuition. A high-mean signal $s_{1}$ induces a bull spread (panel 3b): a long–short vertical spread position that profits from upward revaluation while limiting exposure in far out-of-the-money states. Conversely, a low-mean signal $s_{2}$ induces a bear spread (panel 3d), concentrating downside exposure while economizing on extreme strikes that contribute little to expected profits under the signal-implied distribution.
@@
Left column: the insider’s private signal, indicated by the shaded payoff distributions. Right column: the corresponding informed demands $W^{*}$.
@@
Price Impact. Three states are indicated: $x_{1}$, $x_{2}$, and $x_{3}$.
@@
Suppose the signal selects between two normal payoff distributions with a common mean and variances $\sigma_{1}^{2}>\sigma_{2}^{2}$. Private information therefore concerns the volatility of the payoff rather than its level. In terms of the information flow (29), equilibrium maps this volatility signal into the option portfolios widely used in practice: a straddle when volatility is high, and a butterfly when volatility is low. [^17] See again Hull (2015) for standard definitions of straddles and butterflies.
@@
Portfolio intuition. Under the high-volatility signal $s_{1}$, the distribution places relatively more weight on tail outcomes. Accordingly, the informed demand loads on the wings and avoids the center (panel 4b), producing a straddle-type position—a wing-loaded long-volatility portfolio. Conversely, under the low-volatility signal $s_{2}$ the payoff distribution is more concentrated near the mean. Panel 4d shows that informed demand flips sign across strikes, producing a butterfly position: it concentrates exposure near the center and shorts wing payoffs that are unlikely to realize under low volatility.
@@
Left column: signal-implied payoff distributions (shaded). Right column: corresponding informed demands $W^{*}$.
@@
Suppose the signal selects between two skew-normal payoff distributions that share the same mean and variance but differ in their skewness parameters, with $\alpha_{1}>\alpha_{2}$. Private information therefore concerns asymmetry—how probability mass shifts between left- and right-tail outcomes holding the first two moments fixed. In equilibrium, this skewness signal maps into a standard skew trade: an option portfolio that tilts exposure toward the tail favored by the signal while offsetting exposure to the opposite tail.